2. Below is a payroll sheet for Opel Company for the month of November 2018. The company is allowed a 1% unemployment compensation rate by the state; the federal unemployment tax rate is 0.8% and the maximum for both is $7,000. Assume a 10% federal income tax rate for all employees and a 7.65% FICA tax on employee and employer on a maximum of $117,500. In addition, 1.45% is charged both employer and employee for an employee’s wages in excess of $117,500 per employee. Tax year-end is Dec. 31.

Employee Earnings to Oct. 31 November Earnings Albert $6,700 $850 Benjamin $7,800 $600 Carlson $113,000 $11,200

a. Calculate the company’s Income Tax Withholding, FICA, SUTA, FUTA for November 2018.

b. Prepare the journal entry to record the payment of the November payroll.

c. Prepare the journal entry to record the company’s November payroll tax expenses.

d. Prepare the journal entry to record the payment of the payroll liabilities created in November. Assume that the company pays all payroll liabilities at the end of each month.



Employee Earnings to Oct. 31 November Earnings Albert $6,700 $850 Benjamin $7,800 $600 Carlson $113,000 $11,200
Answer #1
a. Employee Earnings to Oct. 31 Nov earnings Income tax withholding FICA SUTA FUTA
a b c=b*10% d=b*7.65% e
(Note:2)
f
(Note:3)
Albert 6700 850 85 65.025 3 2.4
(7000-6700)*1% (7000-6700)*0.8%
Benjamin 7800 600 60 45.9 0 0
Carlson 113000 11200 1120 441.4 0 0
(Note:1)
Total 1265 552.325 3 2.4
Note:1
FICA limit is $ 117500
Upto $ 117500 - 7.65%
> $ 117500 -1.45%
FICA=(117500-113000)*7.65%+(11200-4500)*1.45%=344.25+97.15=$ 441.40
Note:2
SUTA limit is $ 7000
Upto $ 7000-1%
> $ 7000-Nil
Note:3
FUTA limit is $ 7000
Upto $ 7000-0.8%
> $ 7000-Nil
b. Journal entry:
Date Account titles and explanation Debit Credit
Nov 30. Salaries expense 12650
Income tax withholdings payable 1265
FICA tax payable 552.325
Cash (Plug) 10832.675
(Payment of Nov. payroll)
c. Journal entry:
Date Account titles and explanation Debit Credit
Nov 30. Payroll tax expense 557.725
FICA tax payable 552.325
SUTA payable 3
FUTA payable 2.4
(Payroll tax expenses recorded)
d. Date Account titles and explanation Debit Credit
Nov 30. Income tax withholdings payable 1265
FICA tax payable (552.325+552.325) 1104.65
SUTA payable 3
FUTA payable 2.4
Cash 2375.05
(Payroll liabilities paid)